OpenStudy (anonymous):

3r-2(2r-1)=5(2r-4) my answer is r=2

OpenStudy (angela210793):

3r-4r+2=10r-20 -r+2=10r-20 -11r=-22 r=2 :D

OpenStudy (anonymous):

\[3r-4r+2=10r-20\] \[-r-10r=-20-2\] \[-11r=-22\] \[r=\frac{-22}{-11}\] \[r=2\] u r right good work

OpenStudy (anonymous):

thanks